What is the capacitance of a parallel plate capacitor with metal plates, each of area 1.00m², separated by 2.00mm? What charge is stored in this capacitor if a voltage of 3.00×10³V is applied to it?

1
Expert's answer
2022-03-08T12:19:22-0500

Capacitance

"C=\frac{A\epsilon_0}{d}"

Put value

"C=\frac{1\times8.85\times10^{-12}}{2\times10^{-3}}=4.43n F"

Charge

"Q=CV"


"Q=3\times10^3\times4.43\times10^{-9}=13.27\times10^{-6}C"

"Q=13.27\mu C"